OpenStudy (jiteshmeghwal9):

If 2x+y=3 and xy=1 the value of \(\LARGE{(x+y)^{(x-y)}}\) is :-

hartnn (hartnn):

x(3-2x)=1 and you can easily get 2 values of 'x'

hartnn (hartnn):

x=1,1/2 y=1,2 when x=y=1, that expression =1 when x=1/2, =2, that expression =..... ?

OpenStudy (jiteshmeghwal9):

from where do u gt x(3-2x)=1

hartnn (hartnn):

2x+y=3 and xy=1

hartnn (hartnn):

y=3-2x x(3-2x)=1

OpenStudy (jiteshmeghwal9):

ohh ! okay

hartnn (hartnn):

when x=1/2, y=2, that expression =(5/2)^(-3/2) weird.

OpenStudy (jiteshmeghwal9):

but answer is 1.

hartnn (hartnn):

when x=1,y=1, answer = 1, yes. is there a restriction that x,y are integers .....else you get 2 answers.

OpenStudy (jiteshmeghwal9):

no there is no restriction

OpenStudy (jiteshmeghwal9):

The questio is only asking for the answer of the equation

OpenStudy (jiteshmeghwal9):

It has the following options:- a. 0 b. 2 c.100 d. 1

hartnn (hartnn):

then best choice =1

hartnn (hartnn):

ignore other answer.

OpenStudy (jiteshmeghwal9):

So, d is the answer. Isn't it ??

OpenStudy (jiteshmeghwal9):

ok!

OpenStudy (jiteshmeghwal9):

Thanx !